complete the following proofs within the prescribed
Given: 2x+6=3+5/3x Prove: x=-9

Answer by stanbon(75887)

Given: 2x+6=3+(5/3)x Prove: x=-9
----
Substitute -9 for x to get:
-18+6 = 3 + (5/3)(-9)
-12 = 3 - 15
-12 = -12
----------
Cheers,
Stan H.
--------------